The GDP deflator formula can be used in a variety of ways. Use it to answer the questions that follow. Round final answers to two decimal places, as needed. Real GDP GDP deflator 2012 $2200.00140.00 nominal GDP in 2012 $ Nominal GDP Real GDP $200.00 2011 $1800.00 $1100.00 2010 $700.00 2010 GDP deflator- 2011 GDP deflator = Nominal GDP GDP deflator 2013 $3500.00 160.00 real GDP in 2013s
Answer #1

GDP deflator = (Nominal GDP/ Real GDP )*100

Real GDP = Nominal GDP / (0.01*GDP deflator)

1) NGDP in 2012 = RGDP X GDP deflator /100

NGDP in 2012 = 2200*140/100

NGDP = 3080

2)

Year N GDP R GDP GDP deflator = (N GDP/RGDP)*100
2010 700 200 (700*100/200) = 350
2011 1800 1100 (1800*100/1100)= 163.64

3)

Year N GDP GDP deflator R GDP = (N GDP *100/ GDP deflator)
2013 3500 160 (3500*100/160) = 2187.5